OpenStudy (anonymous):

lim x->1 (sqrt(x) - 1)/ (x - 1)

hartnn (hartnn):

hint : \((x-1)=(\sqrt x)^2-1^2 \) now use the difference of squares \(a^2-b^2=(a+b)(a-b)\)

OpenStudy (anonymous):

how do i do that?

hartnn (hartnn):

\((x-1)=(\sqrt x)^2-1^2=(\sqrt x-1)(\sqrt x+1)\) got this ? what gets cancelled ?

OpenStudy (anonymous):

okay the sqrt(x) - 1 gets cancelled

hartnn (hartnn):

yes, you can directly put x=1 now.

OpenStudy (anonymous):

okay so i substitute x = 1 into sqrt(x) - 1?

hartnn (hartnn):

no....that got cancelled....what remains ?

OpenStudy (anonymous):

what remains is (x - 1)

hartnn (hartnn):

but we wrote x-1 as \((x-1)=(\sqrt x)^2-1^2=(\sqrt x-1)(\sqrt x+1)\)

OpenStudy (anonymous):

ok

hartnn (hartnn):

\(\dfrac{\sqrt x-1}{x-1}=\dfrac{\sqrt{x}-1}{(\sqrt x-1 )(\sqrt x +1)}=\dfrac{1}{\sqrt x+1}\) got this ? now put x=1

OpenStudy (anonymous):

ok i get 1/2

hartnn (hartnn):

i get the same, its correct.

OpenStudy (anonymous):

ok i get 1/2

hartnn (hartnn):

yes, 1/2 is correct.